A bag contains 4 black and 6 red bags and bag B contains 7 black and 3 red balls.A die is thrown.If 1 or 2 appears on it, then bag A is chosen, otherwise bag B. If two balls are drawn at random WITHOUT REPLACEMENT, five the probability of one of them being red and another being black. What is the significance of the word "without replacement" here??

Grade:11

1 Answers

Saurabh Koranglekar
5 years ago
without replacement means you dont put back the ball that is drawn

1/3 is prob for selection of bag A
2/3 for bg B

red from bag A = 6/10

black from B = 7/10

selecting red from A = (1/3)*(6/10):

similarly, proceed for rest cases arrange using set theory and solve
revert if the problem persists